# MuonEq: Balancing Before Orthogonalization with Lightweight Equilibration

### 30 Mar 2026 v1

@@ 4 Experiments
4.1 Main Result
(a) Training Loss
(b) Validation Loss
(c) Wall-clock time
Figure 3: The training and validation loss curves, plotted against both training tokens and wall-clock time on LLaMA-2 130M.
(a) Training Loss
(b) Validation Loss
(c) Wall-clock time
Figure 4: The training and validation loss curves, plotted against both training tokens and wall-clock time on LLaMA-2 350M.
In this section, we evaluate the performance of the proposed MuonEq optimizers on LLM pretraining. All experiments are conducted using 64 Ascend 910C (64GB) NPUs. Detailed experimental settings are provided in Appendix J.
@@ 4.1 Main Result
Figure 3 reports training loss versus tokens, validation loss versus tokens, and validation loss versus wall-clock time for LLaMA2-130M trained to 10.5B tokens. Figure 4 reports the corresponding validation-loss comparison for LLaMA2-350M trained to 21.0B tokens. Table 1 reports the best validation perplexity, peak memory (GB) usage, and training wall-clock time (Second/Step) for these main comparison runs. Across Figure 3, Figure 4, and Table 1, MuonEq consistently improves over Muon.
@@
Optimizer 130M 350M
Perplexity Memory Time Cost Perplexity Memory Time Cost
AdamW Loshchilov and Hutter (2019) 21.35 12.94 0.21 16.25 29.43 0.53
MARS-AdamW Liu et al. (2025b) 20.03 12.99 0.23 15.62 29.47 0.57
Muon Jordan et al. ; Liu et al. (2025a) 20.52 12.52 0.25 15.21 28.18 0.61
Muon-Nes Jordan et al. ; Liu et al. (2025a); Chang et al. (2025) 20.18 12.52 0.25 15.19 28.18 0.61
@@ 4.2 Ablation Study
We ablate the three MuonEq variants RC, R, and C to isolate pre-orthogonalization geometry. All variants use the same optimizer states, scalar hyperparameters, and Newton–Schulz budget; only the pre-orthogonalization diagonal map changes. We take R as the default variant. All ablations run on 4 RTX Pro6000 (96GB) GPUs; details are deferred to Appendix J.2.
The results consistently favor R. Among the three static variants, R performs best on both CIFAR-10/ResNet-18 and FineWeb/GPT2-small, reaching 94.53% test accuracy on CIFAR-10 and 24.94 validation perplexity on FineWeb/GPT2-small. RC is competitive but falls short of R, while C is weaker in this hidden-weight setting. Full static-variant comparisons appear in Tables 3 and 3.
This pattern matches the trade-off in Eq. (5) and the analysis in Section 3. RC provides stronger two-sided spectral correction and can reduce the finite-step Newton–Schulz term more aggressively, but also introduces larger preconditioning bias. For the hidden matrix weights targeted by MuonEq, the more relevant one-sided geometry is row-sided, so R offers the best trade-off between spectral correction and geometric alignment. This interpretation is consistent with (Xu et al., 2026b). Comparisons with (Pethick et al., 2025) and (Glentis et al., 2025) require care because their row/column terminology depends on parameter layout and task setting; we discuss this point in the appendix J.3.
@@
Table 3: Ablation on Fineweb-10.5B tokens with GPT2-small.
@@
AdamW 26.60
Muon-Nes 25.31
@@
5 Related Work
MuonEq is viewed as a lightweight pre-balanced Muon family rather than as a fully whitened optimizer. It inserts diagonal equilibration before Newton–Schulz orthogonalization and includes three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C), with R as the default choice for the hidden matrix weights considered in this paper. All three use only row/column squared-norm statistics and preserve the $\mathcal{O}(m+n)$ state profile.
Muon and orthogonalized-update optimizers. Muon Jordan et al. introduced the core design of orthogonalizing matrix-valued momentum with a small fixed number of Newton–Schulz iterations. Subsequent work has expanded this idea into a broader family of orthogonalized-update optimizers, including post-orthogonalization normalization and rescaling methods such as Muon+, NorMuon, and AdaMuon Zhang et al. (2026b); Li et al. (2025); Si et al. (2025); adaptive stepsize variants such as AdaGO and NAMO Zhang et al. (2025, 2026a); distributed and communication-aware extensions such as Dion and MuLoCo Ahn et al. (2025); Thérien et al. (2025); reduced-state implementations Gupta et al. (2025); and fine-tuning-oriented variants Page et al. (2025). These methods mainly alter the orthogonalized update itself or the surrounding optimization machinery. By contrast, MuonEq modifies the matrix fed into orthogonalization through lightweight pre-orthogonalization equilibration.
Theory of Muon-style methods. Recent theory interprets Muon as steepest descent under spectral-norm or nuclear-norm geometry Bernstein and Newhouse (2024); Chen et al. (2025), establishes stochastic and nonconvex convergence guarantees Shen et al. (2025); Chang et al. (2025); Nagashima and Iiduka (2026); Zhang and Lin (2026), and analyzes practical ingredients such as weight decay, critical batch size, and inexact orthogonalization Sato et al. (2025); Gyu Yeol Kim (2026); Shulgin et al. (2025). This literature is directly relevant to MuonEq because we retain the same finite-step orthogonalization primitive as Muon and change only the geometry of its input. Our analysis further highlights the role of input spectrum in finite-step Newton–Schulz and clarifies how diagonal equilibration trades stronger spectral correction against preprocessing bias.
Structured preconditioning, factorized statistics, and whitening-inspired methods. A separate line of work studies matrix-aware preconditioning, including Shampoo Gupta et al. (2018), K-FAC Martens and Grosse (2015); Eschenhagen et al. (2023), SOAP Vyas et al. (2024), Sketchy Feinberg et al. (2023), and ASGO An et al. (2025). Adafactor Shazeer and Stern (2018) is especially relevant from a systems perspective because it shows how row- and column-factorized second-moment statistics reduce optimizer state to $\mathcal{O}(m+n)$ for matrix-valued parameters. Recent Muon-adjacent whitening or preconditioned-polar methods, including FISMO and Mousse Xu et al. (2026a); Zhang et al. (2026c), pursue a similar high-level goal: improve the geometry before imposing spectral constraints. The key difference is that these methods rely on richer matrix- or Kronecker-structured preconditioners, whereas MuonEq restricts the intervention to diagonal row/column equilibration. In this sense, MuonEq is closer to low-cost equilibration than to full whitening.
Orthogonalization algorithms. There is also a complementary numerical literature on faster and more accurate polar/orthogonalization routines, including improved Newton–Schulz variants and matrix-sign methods such as CANS and Polar Express Grishina et al. (2025); Amsel et al. (2025). This literature matters because Muon-style optimizers operate in the regime of approximate orthogonalization, where only a small number of iterations is affordable; MuonEq is designed explicitly for this setting.
@@ 6 Conclusion
We study how to improve the matrix geometry seen by finite-step orthogonalization in Muon-style optimization without the cost of full whitening. MuonEq does so with a lightweight family of pre-orthogonalization equilibration schemes: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). These variants modify the input to Newton–Schulz orthogonalization while preserving the $\mathcal{O}(m+n)$ memory profile of factorized methods. For the hidden matrix weights studied here, we use R as the default variant.
The analysis shows that finite-step orthogonalization is highly sensitive to the input spectrum and that row/column normalization acts as a zeroth-order whitening surrogate by removing marginal scale mismatch. It also clarifies the trade-off within the MuonEq family: RC gives stronger two-sided spectral correction, whereas R yields cleaner one-sided geometry and the main $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex guarantee. Empirically, the default R variant consistently outperforms Muon in LLaMA2 pretraining on C4 across evaluated model sizes and token budgets. Ablations over RC, R, and C further support R as the strongest default for the hidden-weight setting considered here.
@@
Appendix A Proofs of Theorem 3.1
@@
The claim for $k=0$ is immediate from
@@
Substituting this into the iteration and using $\mathbf{U}^{\top}\mathbf{U}=I_{r}$, we obtain
@@
Moreover, we have
$\operatorname{Orth}(\mathbf{G})=\mathbf{U}\mathbf{V}^{\top}=\mathbf{U}I_{r}\mathbf{V}^{\top},\quad\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})=\mathbf{U}\Bigl(\operatorname{diag}(s_{1}^{(k)},\dots,s_{r}^{(k)})-I_{r}\Bigr)\mathbf{V}^{\top}.$
Since left and right multiplication by matrices with orthonormal columns preserves the Frobenius norm, it follows that
$\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}^{2}=\sum_{i=1}^{r}(s_{i}^{(k)}-1)^{2},$
which proves the exact error representation.
We next prove the lower bound. Define
$q(t):=a+bt+ct^{2},\qquad t\in[0,1].$
Then
$q^{\prime}(t)=b+2ct=-4.7750+4.0630t<0,\qquad t\in[0,1].$
Hence $q$ is strictly decreasing on $[0,1]$, and in particular
$0<q(1)=a+b+c=0.701\leq q(t)\leq q(0)=a,\qquad t\in[0,1].$
Fix any $i\in\{1,\dots,r\}$ and $k\geq 0$. If $a^{k}\sigma_{i}/\alpha\geq 1$, then trivially
$|1-s_{i}^{(k)}|\geq 0=\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
@@
The case $j=0$ is immediate. Suppose the claim holds for some $j<k$. Then $s_{i}^{(j)}\in(0,1)$, so $(s_{i}^{(j)})^{2}\in(0,1)$ and thus
@@
Combining the two cases, we conclude that, for every $i$,
@@
Squaring both sides, summing over $i$, and invoking the exact error representation established above yields
$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}^{2}\right)^{1/2}.$
@@
Remark A.1.
Since
$\phi(1)=a+b+c=0.701\neq 1,$
the scalar map induced by the above Muon coefficients does not have $1$ as a fixed point. Therefore, $T_{\varepsilon}$ in the theorem should be interpreted as the first hitting time of the error threshold $\varepsilon$, rather than as the number of iterations required for asymptotic convergence to $\operatorname{Orth}(\mathbf{G})$.
Appendix B Lemmas for Proposition 3.2
B.1 Lemma B.1
Lemma B.1.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be symmetric positive definite, and let $\mathbf{E}\in\mathbb{R}^{n\times n}$ be symmetric. Define $g(\mathbf{X})=\mathbf{X}^{1/2},f(\mathbf{X})=\mathbf{X}^{-1/2}$. Then $g$ and $f$ are Fréchet differentiable at $\mathbf{A}$. If $\mathbf{S}=\mathbf{A}^{1/2},\mathbf{L}=L_{g}(\mathbf{A};\mathbf{E})$, then $\mathbf{L}$ is the unique symmetric solution of the Sylvester equation
@@
Appendix C Proofs of Proposition 3.2
@@
Apply Lemma B.1 with $\mathbf{A}=\mathbf{D}_{c}^{2},\mathbf{E}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}$. Then
@@
Moreover, by Lemma B.2, we have
@@
Applying Lemma B.1 with $\mathbf{A}=\mathbf{D}_{r}^{2},\mathbf{E}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}$ gives
@@
Again by Lemma B.2, we have
@@
Appendix D Normalization and whitening
This appendix records the normalization–whitening estimates omitted from the main text. We use the notation of Proposition 3.2. All $\mathcal{O}(\cdot)$ terms are understood in operator norm; the implied constants may depend on the fixed matrices in Proposition 3.2, but not on the perturbation terms.
Proposition D.1.
The following perturbative statements hold.
  1. (a)
    In the column/right setting of Proposition 3.2, if $\|\mathbf{C}_{c}\|_{2}$ is small enough, then $\operatorname{Orth}_{r}(\mathbf{N}_{c})=\mathbf{N}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{N}_{c}-\tfrac{1}{2}\mathbf{N}_{c}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$, so $\|\operatorname{Orth}_{r}(\mathbf{N}_{c})-\mathbf{N}_{c}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$. Also,

### 10 May 2026 v2

@@ 4 Experiments
4.1 Main Results
In this section, we evaluate the performance of the proposed MuonEq optimizers on LLM pretraining. We use the same 8-node Ascend 910C cluster for all LLaMA experiments; 130M uses 32 NPUs, whereas 350M and 1B use 64 NPUs. Detailed experimental settings are provided in Appendix K. We will refer to the MuonEq with Nesterov acceleration enabled as MuonEq-Nes.
Refer to captionRefer to captionRefer to caption
((a))
Refer to captionRefer to captionRefer to caption
((b))
Refer to captionRefer to captionRefer to caption
((c))
Figure 3: Comparison of LLaMA2 130M, 350M, and 1B from three perspectives. In each column, the top panel corresponds to the 130M model, the middle panel corresponds to the 350M model, and the bottom panel corresponds to the 1B model. From left to right, the panels show training loss versus tokens, validation loss versus tokens, and validation loss versus wall-clock time.
@@ 4.1 Main Results
Optimizer 130M 350M 1B
Perplexity Memory Time Perplexity Memory Time Perplexity Memory Time
AdamW Loshchilov and Hutter (2019) 21.35 12.94 0.21 16.25 29.43 0.53 14.71 39.56 2.47
MARS-AdamW Yuan et al. (2024) 20.03 12.99 0.23 15.62 29.47 0.57 13.22 39.57 2.50
Muon Jordan et al. ; Liu et al. (2025a) 20.52 12.52 0.25 15.21 28.18 0.61 13.06 34.81 2.76
Muon-Nes Jordan et al. ; Liu et al. (2025a); Chang et al. (2025) 20.18 12.52 0.25 15.19 28.18 0.61 13.24 34.81 2.76
@@ 4.2 Ablation Study
We ablate the three static MuonEq-Nes variants, RC, R, and C, to isolate the effect of pre-orthogonalization geometry. All ablations run on 4 RTX Pro6000 (96GB) GPUs. Detailed experimental settings are provided in Appendix K.2.
Tables 3 and 3 show that, with results averaged over three random seeds, R is consistently the strongest static variant across both benchmarks, achieving higher test accuracy than Muon-Nes on CIFAR-10 with ResNet-18 and lower validation perplexity on FineWeb-10.5B tokens with GPT2-small.
((a))
((b))
((c))
((d))
Figure 4: (a,b) Validation perplexity and step time vs. Newton–Schulz iterations $K$ on GPT2-small/FineWeb (10.5B tokens); (c,d) validation perplexity heatmaps over learning rate and momentum at $K=5$ (2.6B tokens).
@@
Table 3: Ablation on FineWeb with GPT2-small.
@@
AdamW 26.48±0.12
Muon-Nes 25.23±0.08
AdaMuon 24.99±0.13
Muon+ 25.12±0.08
Mousse 25.16±0.10
@@
Figure LABEL:fig:muon-val and LABEL:fig:muon-speed show that the default R variant of MuonEq-Nes is markedly less sensitive to the NS iteration count $K$ than Muon-Nes: its validation perplexity remains stable over a wider range of $K$, while step time stays comparable throughout the sweep. Complementary robustness evidence is given in Figure LABEL:fig:heatmap-muon-k5 and LABEL:fig:heatmap-muoneq-k5, where MuonEq-Nes (R) exhibits a broader low-perplexity region over learning-rate and momentum sweeps than Muon-Nes, especially at $K=5$. This pattern is consistent with Eq. (5) and Section 3. RC applies stronger two-sided spectral correction and can reduce finite-step NS error more aggressively, but at the cost of larger preconditioning bias. For the hidden-weight settings we tested, R provides a favorable trade-off between spectral correction and preprocessing bias, consistent with (Xu et al., 2026b). Comparisons with (Pethick et al., 2025) and (Glentis et al., 2025) require caution, as row/column terms vary by layout and task; see Appendix K.3.
@@ 5 Conclusion
We improve the matrix geometry encountered by finite-step orthogonalization in Muon-style optimization without resorting to full whitening. MuonEq reshapes the input to Newton–Schulz using lightweight pre-orthogonalization equilibration in three forms, RC, R, and C; R is used by default for hidden weights. Theoretically, these schemes serve as zeroth-order whitening surrogates: RC provides stronger two-sided correction, whereas R yields cleaner one-sided geometry and supports the main $\widetilde{\mathcal{O}}(T^{-1/4})$ guarantee. Empirically, R consistently outperforms Muon in LLaMA2 pretraining on C4 across all tested scales and budgets.
@@
Appendix A Related Work
MuonEq is viewed as a lightweight pre-balanced Muon family rather than as a fully whitened optimizer. It inserts diagonal equilibration before Newton–Schulz orthogonalization and includes three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C), with R as the default choice for the hidden matrix weights considered in this paper. All three compute row/column squared norms from the current momentum and add no persistent optimizer state beyond Muon.
Muon and orthogonalized-update optimizers. Muon [22] introduced the core design of orthogonalizing matrix-valued momentum with a small fixed number of Newton–Schulz iterations. Subsequent work has expanded this idea into a broader family of orthogonalized-update optimizers, including post-orthogonalization normalization and rescaling methods such as Muon+, NorMuon, and AdaMuon [59, 28, 48]; adaptive stepsize variants such as AdaGO and NAMO [57, 58]; distributed and communication-aware extensions such as Dion and MuLoCo [1, 49]; reduced-state implementations [16]; and fine-tuning-oriented variants [34]. These methods mainly alter the orthogonalized update itself or the surrounding optimization machinery. By contrast, MuonEq modifies the matrix fed into orthogonalization through lightweight pre-orthogonalization equilibration.
Theory of Muon-style methods. Recent theory interprets Muon as steepest descent under spectral-norm or nuclear-norm geometry [4, 8], establishes stochastic and nonconvex convergence guarantees [46, 6, 33, 61], and analyzes practical ingredients such as weight decay, critical batch size, and inexact orthogonalization [42, 18, 47]. This literature is directly relevant to MuonEq because we retain the same finite-step orthogonalization primitive as Muon and change only the geometry of its input. Our analysis further highlights the role of input spectrum in finite-step Newton–Schulz and clarifies how diagonal equilibration trades stronger spectral correction against preprocessing bias.
Structured preconditioning, factorized statistics, and whitening-inspired methods. A separate line of work studies matrix-aware preconditioning, including Shampoo [17], K-FAC [32, 11], SOAP [51], Sketchy [13], and ASGO [3]. Adafactor [45] is especially relevant from a systems perspective because it shows how row- and column-factorized second-moment statistics reduce optimizer state to $\mathcal{O}(m+n)$ for matrix-valued parameters. Recent Muon-adjacent whitening or preconditioned-polar methods, including FISMO and Mousse [54, 60], pursue a similar high-level goal: improve the geometry before imposing spectral constraints. The key difference is that these methods rely on richer matrix- or Kronecker-structured preconditioners, whereas MuonEq restricts the intervention to diagonal row/column equilibration. In this sense, MuonEq is closer to low-cost equilibration than to full whitening.
Orthogonalization algorithms. There is also a complementary numerical literature on faster and more accurate polar/orthogonalization routines, including improved Newton–Schulz variants and matrix-sign methods such as CANS and Polar Express [15, 2]. This literature matters because Muon-style optimizers operate in the regime of approximate orthogonalization, where only a small number of iterations is affordable; MuonEq is designed explicitly for this setting.
Appendix B Proofs of Theorem 3.1
@@ Appendix A Related Work
Let $\alpha>0$. For $k=0$, the desired form follows directly from
@@
Inserting this expression into the iteration and using $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{r}$ gives
@@
We also have
$\operatorname{Orth}(\mathbf{G})=\mathbf{U}\mathbf{V}^{\top}=\mathbf{U}\mathbf{I}_{r}\mathbf{V}^{\top},\quad\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})=\mathbf{U}\Bigl(\operatorname{diag}(s_{1}^{(k)},\dots,s_{r}^{(k)})-\mathbf{I}_{r}\Bigr)\mathbf{V}^{\top}.$
Left and right multiplication by matrices with orthonormal columns preserves the Frobenius norm, so
$\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}^{2}=\sum_{i=1}^{r}(s_{i}^{(k)}-1)^{2}.$
This proves the exact error representation.
We now establish the lower bound. Recall that
$q(t):=a+bt+ct^{2}.$
By assumption,
$0<q(t)\leq a,\qquad t\in[0,1].$
Fix $i\in[r]$ and $k\geq 0$. If $a^{k}\frac{\sigma_{i}}{\alpha}\geq 1,$ then
$\bigl|1-s_{i}^{(k)}\bigr|\geq 0=\left(1-a^{k}\frac{\sigma_{i}}{\alpha}\right)_{+}.$
@@
The statement is immediate for $j=0$. Suppose it holds for some $j<k$. Then $s_{i}^{(j)}\in(0,1)$, and hence $(s_{i}^{(j)})^{2}\in(0,1)$. Therefore,
@@
Combining the two cases gives, for every $i$,
@@
Finally, set $\alpha=\|\mathbf{G}\|_{F}$. Squaring both sides, summing over $i$, and using the exact error representation together with
$\frac{\sigma_{i}}{\|\mathbf{G}\|_{F}}=\frac{\sigma_{i}/\sigma_{1}}{\|\mathbf{G}\|_{F}/\sigma_{1}}=\frac{1}{\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}},$
we obtain
$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\left(1-\frac{a^{k}}{\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}}\right)_{+}^{2}\right)^{1/2}.$
Moreover,
$\tau_{i}:=\log_{a}\frac{\|\mathbf{G}\|_{F}}{\sigma_{i}}=\log_{a}\left(\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}\right),$
so
$\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\qquad\tau_{r}=\log_{a}\left(\kappa(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\right),\qquad\tau_{r}-\tau_{1}=\log_{a}\kappa(\mathbf{G}).$
@@
Appendix C Lemmas for Proposition 3.2
C.1 Lemma C.1
Lemma C.1.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be symmetric positive definite, and let $\mathbf{E}\in\mathbb{R}^{n\times n}$ be symmetric. Define $g(\mathbf{X})=\mathbf{X}^{1/2},f(\mathbf{X})=\mathbf{X}^{-1/2}$. Then $g$ and $f$ are Fréchet differentiable at $\mathbf{A}$. We use $L_{h}(\mathbf{A};\mathbf{E})$ to denote the Fréchet derivative of a matrix function $h$ at $\mathbf{A}$ applied to $\mathbf{E}$, i.e., $h(\mathbf{A}+\mathbf{E})=h(\mathbf{A})+L_{h}(\mathbf{A};\mathbf{E})+o(\|\mathbf{E}\|_{2})$. If $\mathbf{S}=\mathbf{A}^{1/2},\mathbf{L}=L_{g}(\mathbf{A};\mathbf{E})$, then $\mathbf{L}$ is the unique symmetric solution of the Sylvester equation
@@
Appendix D Proofs of Proposition 3.2
@@
Apply Lemma C.1 with $\mathbf{A}=\mathbf{D}_{c}^{2},\mathbf{E}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}$. Then
@@
Moreover, by Lemma C.2, we have
@@
Applying Lemma C.1 with $\mathbf{A}=\mathbf{D}_{r}^{2},\mathbf{E}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}$ gives
@@
Again by Lemma C.2, we have
@@
Appendix E Normalization and whitening
This appendix collects the normalization–whitening estimates that are omitted from the main text. We use the notation of Proposition 3.2. All $\mathcal{O}(\cdot)$ terms are taken in operator norm; the implied constants may depend on the fixed matrices in Proposition 3.2, but not on the perturbation terms.
Proposition E.1.
The following estimates hold.
In the column/right setting of Proposition 3.2, if $\|\mathbf{C}_{c}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{r}(\mathbf{N}_{c})=\mathbf{N}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{N}_{c}-\tfrac{1}{2}\mathbf{N}_{c}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{r}(\mathbf{N}_{c})-\mathbf{N}_{c}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$. Moreover,